From acd@roe.ac.uk Fri Jul 4 16:37:04 2014 Date: Thu, 3 Jul 2014 15:06:39 +0000 From: Clive Davenhall To: Bob Mann , Andy Lawrence , "Read, Mike" , Eckhard Sutorius , Mark S. Holliman , Nicholas Cross , Nigel C. Hambly , Rob Blake , Ross Collins , Stelios Voutsinas , Dave Morris , Clive Davenhall Cc: "dante@astro.puc.cl" , James Dunlop , "j.p.emerson@qmul.ac.uk" , ktn keithnoddle.org , "Lorenzo.Rimoldini@unige.ch" , "mcioni@aip.de" , "matt.jarvis@astro.ox.ac.uk" , "mike@ast.cam.ac.uk" , "nigel.metcalfe@durham.ac.uk" , "norman@astro.gla.ac.uk" , "pwl@star.herts.ac.uk" , "rgm@ast.cam.ac.uk" , "Simon.Dye@nottingham.ac.uk" , "s.j.warren@ic.ac.uk" , "tkerr@jach.hawaii.edu" , "tom.shanks@durham.ac.uk" Subject: WFAU VDFS Science Archives progress meeting minutes, 20 June 2014 WFAU VDFS Science Archives progress meeting minutes, 20 June 2014 Present: NJC, RPB, ETWS, RGM, MSH, SV, ACD Apologies: MAR, RSC, NCH, KTN, AL DoNM: 10am Friday 4 July 2014 in the Villas meeting room, MAR in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: ACD, RPB to formulate a policy for backups and for checking that data could be recovered from the backups. Done. A policy had been formulated but some questions remain, which resulted in the following two new actions: *NEW ACTION: RPB to write a script to backup the SVN repository. *NEW ACTION: RPB estimate the number of additional tapes that would be need to backup all the data on disk. ACTION: RPB to upgrade STARLINK on curation servers. Done. Most systems had been upgraded but, following discussion, the Web servers had not been. Actions partly discharged but continuing: ---------------------------------------- None. Actions carried forward from 9/5/2014 meeting: ---------------------------------------------- ACTION: RPB to put all prices of additional items associated with the backup policy on the Twiki pages. Continues. ACTION: RGM to investigate the size and content (how does it differ from DR9?) of DR10 to see if it's worth us hosting and whether it is necessary to keep all of the other releases. Continues. RGM had started a wiki page for WFAU SDSS strategy http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan). ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. ACTION: MAR to document issue with ATLAS unpaired sources and give examples of recovering matches. Continues. MAR was not present. ACTION: ETWS to investigate modifying post CU4 ops to calculate and store illumination correction values. Continues. A script had been written but needs to be tested on real data, ACTION: ACD to arrange follow-up VVV brainstorming meeting. Continues. RSC was organising the testing and the meeting should be organised ICW him. Actions carried forward from 23/5/2014 meeting: ---------------------------------------------- ACTION: RPB to create a wiki page to discuss ways of generating unique objIDs. Continues. A draft wiki page was available and comments on it had been requested. ACTION: RPB/ACD/MSH to see if a new database node could be ordered. Continues. ACD needed to send an approval of the purchase to Nathalie Dupin. Specific points and new actions: ------------------------------- Project management: * Nothing to report. WFCAM, VISTA and VST updates: * The telescopes were operating normally. * The ordering of tapes to permit the archiving or raw WFCAM data was on-track. * NJC noted that Janet Drew had made a request that would require a lot of work and could involve ingesting additional data. Discussions were still in progress. ACTION: NJC to circulate the relevant e-mail messages. Comments and issues arising from CASU minutes: * No minutes had been received since the previous meeting. Review of Deadlines: * NJC had negotiated deadlines for the forthcoming ESO releases (and the dates on the Web page would need updating): VIDEO - September, VIKING - late June or early July. ACTION NJC: update ESO VISTA releases on the Web page. * ETWS noted that the July OSA deadline would not be met because of a lack of disk space. Networking: * Nothing to report. WSA/VSA/OSA/GES Operations: * RPB commented on the VIKING and VMC release databases: the VMC detection table was empty. This problem was unsolved. * NJC reported: - VIKING: was progressing slowly, - VMC: there was new detection software and tables of eclipsing binaries, - VVV almost ready to go. * ETWS noted that there was nothing to report on the OSA. * ACD noted that work continued on preparation of the GES iDR2 release, it was very late and problems continued to be found. Hardware and Systems: * MSH noted that he and the ATC computing support had been warned by someone in Geosciences that they had been approached by an individual selling tapes that were claimed to come from the ROE. MSH suggested conducting a tape inventory to check that none are missing. ACTION: MSH, RPB conduct a tape inventory to check if any tapes are missing. * MSH also noted that purchase of the new database node was progressing. Software: * NJC noted that he had made some small changes to the VISTA software for the forthcoming ESO releases. The changes were mainly for the VVV where the code to produce the DR2 synoptic source list had been simplified. He noted that the new VMC data were well documented. * ETWS reported that he had written/updated the code to ingest the uncorrected catalogues as well as the CASU list-driven photometry catalogues into the OSA. This work included updating the schema files to include the new *DetectionUncorr* tables as well as including the new attributes for the APASS zero points. * SV noted that the infrastructure to test the new ADQL query interface to the OSA was in place and he planned to use it to run previously-submitted queries through the new interface. He also noted that a 'raw SQL' option had been added to the ADQL interface. Survey Data Release: * RPB noted that VIKING and VMC releases were imminent. Non-survey Data Release: * None; there had been insufficient time to work on these and they were low priority. ACTION: MSH produce a costed list of the additional kit necessary to permit reprocessing. ACTION: ACD, RGM: review finances. * MSH noted that a 4 Tbyte SSD had recently been announced and was potentially very useful. Astrogrid deployment, VO & Data Analysis services: * SV noted that at the recent IVOA meeting good progress had been reported on the additional syntax for ADQL, which contained features that would be useful for the new OSA interface. * MSH noted that the old AstroGrid registry would become redundant within the next 9 months. In that time TOPCAT will be using a different registry, the EURO-VO registry has been upgraded and RegTap implementations are becoming available within registries. Miscellaneous: * Nothing to report. -- Scanned by iCritical.